OpenStudy (missingno):

Which of the following is not a way to represent the solution of the inequality 2(x − 1) greater than or equal to −12? x greater than or equal to −5 x less than or equal to −5 −5 less than or greater to x A number line with a closed circle on negative 5 and shading to the right

OpenStudy (will.h):

you were given the following inequality \[2(x - 1) \ge -12\] you should solve for x so that you find the correct option. now the 1st thing to do here is to use distrubtive property in the left side so we should multiply 2 * x and x * -1 which would give us \[2x - 2 \ge -12\] now we need to isolate x. 1st we should get rid of -2 by adding it to both sides. \[2x \ge -10\] And now divide both sides on 2 \[x \ge -5\]

OpenStudy (will.h):

so that gives you the answer.

OpenStudy (missingno):

thanks

OpenStudy (will.h):

you welcome :)